Ionization constant of CH3COOH is 1.8 x 10-5 . The concentration of H+ ion in 0.1M solution is?

Grade:

1 Answers

Arun
5 years ago
H+] ^2 = 0.1*1.8*10^-5 = 1.8*10^-6
So [H+] = 1.342*10^-3
Hope it helps

Regards
Arun (askIITians forum expert)